# An Adaptive Computational Intelligence Approach to Personalised Health and Immune Age Characterisation from Common Hematological Markers

### v5

@@ Abstract
We introduce a simulated digital model that learns a person’s optimal blood health over time. Using a learning adaptive algorithm, our model provides a risk assessment score that compares an individual’s chronological age from birth to an estimation of a biological immune age derived from the score. We demonstrate its efficacy against real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from 100K data records in the Centers for Disease Control and Prevention’s National Health and Nutrition Examination Survey (CDC NHANES) that spans 13 years. We find that the score is informative when distinguishing healthy individuals from those with diseases, both self-reported and abnormal blood tests manifested, providing an entry-level score for patient triaging. The main contributions of this work is that the risk assessment score is not a machine learning black-box approach but a transparent adaptive learning algorithm and its calculation only requires existing common markers from a standard and most popular medical (and blood) test widely performed worldwide. This may allow a massively scalable approach to deploying a fast, personalised, sensitive and predictive derivative index without the need or application of a new test, assay, or prospective sampling as opposed to other biological ageing-related scores through a clinical platform and SaMD model.
Keywords: adaptive health learning, personalised health, predictive medicine, twin digital health, blood risk score, risk stratification, population assessment, individual assessment, immune age, biological age, senescence, longevity
@@ 1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores used for initial patient classification or diagnosis and triaging (i.e., assessment of priority for healthcare services) are often qualitative and inter-subjective. Robust quantitative tools with clinical relevance are required to optimise health monitoring choices across all levels of healthcare systems.
@@
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of health care. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [23, 31], clinical acuity [10, 33], and long-term outcomes [30, 32]. For example, the Intermountain risk score [19] provides a complete/full blood count (CBC/FBC) risk score shown to be associated with bleeding and in-hospital mortality based on red blood cell size, originally used as an anaemia predictor [26], and is often reported in FBC/CBC panels.
FBC/CBCs have been found to contain enough information to produce clinically useful scores [3, 18]. At the same time, it has also been shown that although a CBC is a low-cost laboratory test that is almost universally used, its risk predictive information content is often underused [29].
@@
Regular blood testing has been associated with short-term mortality and with the ability to significantly improve emergency department triage [22]. Most risk scores depend on multiple factors external to a single test result in itself and have proven useful when available, even in the context of the immune system [34]. A number of these scores are based on such factors as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits. While numerical scores and colour-based scores are relatively common, to the authors’ knowledge no single general numerical and colour-based blood test risk assessment score exists that is based on a combination of distances from population or personalised average values, in particular for CBC and immune-related data.
In medical practice, a simple risk score based solely on blood test results and agnostic with respect to a specific condition while still being able to quantify the results of a blood test has not been proposed, to the authors’ knowledge. Here we introduce a general numerical and colour-based risk stratification score to quantify any number of abnormal blood markers. The risk model introduced is not data expensive when run ab initio, and yet it is shown to be informative, and its implementation is easily interpretable, while allowing for adaptation to enhance sophistication.
@@
2 Methods
Here we introduce a family of indices based on a distance metric and closed learning algorithm for detecting rapid, small (precise), and personalised changes in blood test results. The indexes have the advantage of being fully explainable, unlike those featured in other statistical or machine-learning approaches. This metric takes the form of a quantitative measure or score that aggregates data from a set of numeric blood test results, combining and condensing them into a single numerical value that allows rapid triage, sorting, and immune age assessment. The main purpose of the score is to learn from, monitor, and provide a quick assessment of divergence from population and personalised reference values.
We will call this risk assessment score ”immune score”, given its connection to the immune system through blood cell subpopulation counts, in particular as registered in the most popular blood test, the FBC or the CBC.
The immune score constitutes a dimensional reduction technique based on a single real-value number and a colour scheme that takes a multidimensional blood test space with dimension size equal to the number of analytes, where each value of a marker or analyte corresponds to the coordinate of that value in that dimension. The score itself can be seen as the norm of a suitable vector transformation that pinpoints the health status of the patient in that space, relative to the test and the set of markers or analytes. Thus, the numerical vector value integrates all the analyte’s space dimensions.
@@ 2.1 Numerical score
This first formulation of the score takes the form of a linear function because it does not take into account possible interactions between analytes. CBC blood cell counts can interact through immune system responses. For example, an infection may increase neutrophils and decrease lymphocytes. RBC parameters can also be impacted by complex environmental cues such as nutrient levels, iron deficiency, stress conditions, or disorders. We also assume that each analyte contributes equally to the immune score (this is likely to be only partially true, as some of the analytes are more or less medically informative than others, with this degree of informativeness itself varying with different conditions. Moreover, not all analytes are independent, and some analytes may be statistically dependent on others—suitable refinements will be introduced later). Some examples of analytes that are independent of each other are RBC and WBC count (they do not directly influence the counts of each other). The counts of some analytes are dependent from each other. Examples include Neutrophils and WBC. In essence, neutrophils are a type of leukocyte (WBC), so the neutrophil count is included in and contributes to the total WBC. Another example is Hemoglobin and hematocrit. Hemoglobin measures the amount of hemoglobin protein in red blood cells. Hematocrit measures the percentage of blood volume taken up by red blood cells. Thus, Hematocrit is mathematically dependent on the hemoglobin concentration because one gives a subset or ratio of the other.
For ease of comparison between successive score versions, all values are normalised within the range $[0,10]$. Following user feedback, we then reverse the scale by subtracting the normalised value from 10. This means that each analyte will contribute with a maximum weight of
@@
The adoption of a score within a maximum range of $[0,10]$ can be approached in different ways. One is to assume no theoretical maximum values for analytes and asymptotically approach the minimum score of 0, without ever reaching it in individual cases. We adopted an alternative approach, capping values to a pre-established “reasonable” maximum, beyond which all specific values mean the same. The pre-established maximum is two standard deviations beyond normal (healthy) ranges, which sets the statistical significance critical value alpha at 0.05 (i.e., 95% confidence interval).
@@
2.2 Colour-based score
@@
Clearly, the CCIS is not a numerical value, but a colour-coded output, like the individual alerts for separate analytes. This clearly distinguishes the CCIS from the normalised immune score.
@@
3 A numerical score derived from the CCIS
We explored an alternative score that would follow the colour-coded scheme, but would also produce a numerical value. The intention was to combine the accessibility of the CCIS and the finer precision of the NIS in a single value. As with the latter, the value would range from 0 to 10.
@@ 3 A numerical score derived from the CCIS
Table 2: Mean values and standard deviations found in the NHANES 2003–2016 database for the 13 analytes of the immune scores (same units as above).
@@
Total White Cells 6.92 3.12 7.01 2.09
@@
Neutrophils 4 1.85 4.16 1.66
Lymphocytes 2.11 2.21 2.11 0.79
Monocytes 0.57 0.21 0.52 0.18
Eosinophils 0.20 0.16 0.17 0.14
Basophils 0.04 0.06 0.04 0.05
The behaviour of Fig. 3 illustrates how values at a remove from healthy values are pushed towards higher values by design, as an alerting mechanism.
3.1 Calculating CCIS and NCCIS
@@
  • $w_{a}=\sqrt{100/13}$ is the maximum normalised distance for each analyte in the normalised vector of a patient’s readings.
  • $e(g,i)$ is the mean (or expected) value for analyte $i$ in group $g$.
  • $\sigma(g,i)$ is the standard deviation from the mean.
  • $n(g,i)=(r_{u}(g,i)-r_{l}(g,i))/2$ is the normal (healthy reference) interval length from the mean value (assuming upper and lower limits are equidistant from the mean, an assumption to be revised in the future, both in the NIS and here in the NCCIS).
@@
For global values the calculation has to take into account both the definition of the CCIS and the fact that each analyte can have proportionally different normal ranges and standard deviations:
@@ 3.2 NHANES Age Prediction
The zero counts were removed from the NHANES data prior to all analysis. However, the sex was replaced by 1 for males and 2 to denote females, for the quantitative analysis. The sex groups were separated by health status as was determined by binning the immune scores and ages by the healthy range thresholds as indicated on Table 1, which are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and for the multivariate measure, the healthy ranges were as follows: normal Mean Corpuscular Hemoglobin (MCH) (27-33), Hematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above, or below these healthy count ranges, the individuals were binned into the unhealthy groups. Further, the error rates analysis for the age prediction between the actual ages and immune-score computed ages were performed using these binned subsets.
Figure 2: Immune score space: The $x$ axis represents the deviation of all the analytes from the expected value, while the $y$ axis measures the total number of analytes. A negative standard deviation indicates that the value is lower than expected and a positive one indicates a larger than expected value. The cut-off seen on the negative $x$ axis shows the fact that real-life values for analytes have a strict lower bound but, technically, have no upper bounds. For instance, it is not possible to have a negative white blood cell count.
Figure 3: The behaviour of NCCIS over synthetically generated patients with differing values and a deviation from the norm with respect to the respective expected values. The $x$ axis measures the stated analyte, while the $y$ axis measures the total number of analytes. In contrast to the behaviour of the immune score (Fig. 2), NCCISS is a non-linear function that displays a step-like behaviour by design, while remaining independent of the number of analytes.
Comparison Sample Size (n) p-value
Healthy vs Unhealthy 7007 vs 6074 P ¡ 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P ¡ 0.0001
Healthy vs Anemia 7007 vs 980 P ¡ 0.0001
Healthy vs HIV 7007 vs 62 P ¡ 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P ¡ 0.0001
Healthy vs Leukemia 7007 vs 18 P ¡ 0.0001
Healthy vs Blood Cancer 7007 vs 8 P ¡ 0.0001
Table 4: Differences in NIS Mean Values. All comparisons were performed using ANOVA.
Figure 4: Distribution of the Normalised Immune Score (NIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2014 database. The Healthy cohort is defined as those individuals with all analytes within the normal ranges as defined by the NHS (1), others comprise the Unhealthy set. The Self rep. healthy cohort compromises individuals that self-reported health of 3 (”Good”) or less The number of individuals in each cohort are shown in table 4.
Comparison Sample Size (n) p-value
Healthy vs Unhealthy 7007 vs 6074 P ¡ 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P ¡ 0.0001
Healthy vs Anemia 7007 vs 980 P ¡ 0.0001
Healthy vs HIV 7007 vs 62 P ¡ 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P ¡ 0.0001
Healthy vs Leukemia 7007 vs 18 P ¡ 0.0001
Healthy vs Blood Cancer 7007 vs 8 P ¡ 0.0001
Table 5: Differences in NCCIS Mean Values. All comparisons were performed using ANOVA.
Figure 5: Distribution of the Normalised Colour-Coded Immune Score (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (1) and Unhealthy otherwise. Self rep. healthy compromises individuals that self-reported health of 3 (“Good”) or less. The number of individuals in eah cohort is shown in table 5. One explanation for the bimodality of some of these plots is the nature of the source data as self-reported where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS. In either case, both score-separating capabilities are highly statistically significant.

### v7

# A Neuro-Symbolic AI Approach to Personal Health Risk Assessment and Immune Age Characterisation using Common Blood Markers

@@ Abstract
We introduce a simulated digital model that learns a person’s baseline blood health over time. Using an adaptive learning algorithm, the model provides a risk assessment score that compares an individual’s chronological age with an estimation of biological age based on common immune-relevant markers used in current clinical practice. We demonstrate its efficacy on real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from 100K data records in the Centers for Disease Control and Prevention’s National Health and Nutrition Examination Survey (CDC NHANES) that spans 13 years. We find that the score is informative in distinguishing healthy individuals from those with diseases, both self-reported and as manifested via abnormal blood test results, providing an entry-level score for patient triaging. The risk assessment score is not a machine learning black-box approach but can interact with ML and DL approaches to help guide, control the attention given to specific features, and assign proper explainable weight to an otherwise transparent adaptive learning algorithm. This approach may allow fast and scalable deployment to personalised, sensitive, and predictive derivative indexes within digital medicine, without the need for a new test, assay, or prospective sampling, unlike other biological ageing-related scores and methods. It demonstrates the potential of clinical informatics and deep medicine in digital healthcare as drivers of innovation in preventive patient care.
Keywords: digital medicine, deep medicine, adaptive health learning, precision healthcare, predictive medicine, digital blood twin, risk stratification score, immune age, senescence, longevity, ageing, deep learning.
@@ 1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores used for initial patient classification or diagnosis and triaging (i.e., assessment of priority for healthcare services) are often qualitative and inter-subjective. Robust quantitative tools with clinical relevance are required to optimise health monitoring choices across all levels of healthcare systems. Digital medicine aims to improve the quality of patient care and outcome trajectories through clinical informatics such as risk assessment metrics grounded in clinical data.
@@
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes in precision medicine. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of healthcare. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [27, 35], clinical acuity [12, 37], and long-term outcomes [34, 36]. For example, the Intermountain risk score [24] provides a complete/full blood count (CBC/FBC) risk score based on red blood cell size, shown to be associated with bleeding and in-hospital mortality. Originally used as an anaemia predictor [30], it is often reported in FBC/CBC panels. As precision healthcare and digital medicine continue to advance, these risk scores will likely integrate more advanced, data-driven approaches to improve prediction accuracy and ensure relevance across a broad spectrum of clinical conditions.
On the one hand, immune age-related scores based on inflammation markers have traditionally required highly specialised and therefore uncommon and expensive tests unsuitable for mass adoption or mass deployment in current clinical pathways, as they are difficult to scale to cover large populations [1, 14, 19, 3, 20]. On the other hand, the most popular medical test, FBC/CBC, has been found to often contain enough information to yield clinically useful insights [5, 23], and even specific disease signatures [29]. While it is a low-cost popular laboratory test that is almost universally used, its risk predictive information content is underused [33], often in favour of more sophisticated omics tests that are also expensive and not currently scalable.
@@
Regular blood testing has been associated with prediction of short-term mortality and with the ability to significantly improve emergency department triage [26]. Most risk scores depend on multiple factors external to a single test result in and of itself, and have proven useful when available, even in the context of the immune system [38]. A number of these scores are based on such factors as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits. While numerical scores and colour-based scores are relatively common, no single general numerical and colour-based blood test risk assessment score exists that is based on a combination of distances from population and personalised average values, in particular for CBC/FBC and immune-related data, used to guide Machine Learning approaches for discrimination of individual health predictors.
Here, we introduce a computational intelligence approach based on or guided by explainable risk stratification scores to quantify any number of blood markers relevant to health and immune age. The risk model introduced is not data expensive when run ab initio, and yet it is shown to be informative and predictive, and implementation remains interpretable even in combination with machine and deep learning, while allowing for clinical adaptation based on patient profiles.
Figure 1: Population reference values and population average inadequacies for personal and precision healthcare: Distributions within ‘healthy’ reference values are heavily skewed or display long tails in most cases, not following normal or uniform distributions. This and current clinical evidence [8, 7] calls for a personalised approach to reference values for abnormal results rather than one-size-fits-all current population approaches to blood test interpretation. The list of analyte labels is in the Supplementary Information.
2 Methods
A complete blood count (CBC) or Full Blood Count (FBC) is the most common clinical blood test that often includes a 5-part differential comprising: the total number of red blood cells (RBCs); the total number of white blood cells (WBCs) and platelets (PLT); haemoglobin (HGB); haematocrit (HCT); estimations related to concentration, shape and volume of red cells (MCV, RDW, MCH and MCHC) and counts of different WBC subtypes, including Lymphocytes (LYMP), Monocytes (MONO), Neutrophils (NEUT), Eosinophils (EOS) and Basophils (BAS). A CBC is a routine blood test used to assess overall health and detect a wide range of disorders, including anaemia, infection and leukaemia.
@@ 2 Methods
We first introduce a family of indices based on a distance metric and closed learning algorithm for detecting rapid, small (precise), and personalised changes in blood test results. These indices have the advantage of being fully explainable, unlike those featured in other, purely statistical or machine-learning approaches.
This metric takes the form of a quantitative measure or score that aggregates data from a set of numerical blood test results, combining and condensing them into a single numerical value that facilitates rapid triage, sorting, and immune age assessment. The main purpose of the score is to learn from, monitor, and provide a quick assessment of divergence from population and personalised reference values.
We will call this risk assessment score the “immune score”, given its connection to the immune system through blood cell subpopulation counts, specifically as registered in the most popular blood test, the FBC or CBC. These scores will then be combined with more traditional Machine Learning (ML) and Deep Learning (DL) approaches, where the scores will help guide the search for specific features, in what constitutes a neuro-symbolic approach to health risk stratification and immune age characterisation in digital medicine.
The immune score constitutes a dimensional reduction technique based on a single real-value number and a colour scheme that takes a multidimensional blood test space with dimension size equal to the number of analytes, where each value of a marker or analyte corresponds to the coordinate of that value in that dimension. The score itself can be seen as the norm of a suitable vector transformation that pinpoints the health status of the patient in that space, relative to the test and the set of markers or analytes. Thus, the numerical vector value integrates all the analyte’s spatial dimensions.
@@ 2.1 Numerical Score
This first formulation of the score takes the form of a linear function because it does not take into account possible interactions between analytes. CBC blood cell counts can be impacted by interactions reflecting immune system responses. For example, an infection may increase neutrophils and decrease lymphocytes. RBC parameters can also be impacted by complex environmental cues such as nutrient levels, iron deficiency, stress conditions, or disorders. We also assume that each analyte contributes equally to the immune score (this is likely to be only partially true, as some of the analytes are more or less medically informative than others, with this degree of informativeness itself varying with different conditions. Moreover, not all analytes are independent, and some analytes may be statistically dependent on others—suitable refinements will be introduced later). Some examples of analytes that are independent of each other are RBC and WBC count (each does not directly influence the other count). The counts of some analytes are dependent on each other. Examples include Neutrophils and WBC. In essence, neutrophils are a type of leukocyte (WBC), so the neutrophil count is included in and contributes to the total WBC. Another example involves Haemoglobin and Haematocrit levels. Haemoglobin measures the amount of haemoglobin protein in red blood cells, while Haematocrit measures the percentage of blood volume taken up by red blood cells. Thus, Haematocrit is mathematically dependent on the haemoglobin concentration because one represents a subset or ratio of the other.
For ease of comparison between successive versions of the score, all values are normalised within the range $[0,10]$. Following user feedback, we then reverse the scale by subtracting the normalised value from 10. This means that each analyte will contribute with a maximum weight of
@@
The adoption of a score within a maximum range of $[0,10]$ can be approached in different ways. One is to assume no theoretical maximum values for analytes and asymptotically approach the minimum score of 0, without ever reaching it in individual cases. We adopted an alternative approach, capping values to a pre-established “reasonable” maximum, beyond which all specific values mean the same. The pre-established maximum is two standard deviations beyond normal (healthy) ranges, which sets the statistically significant critical value alpha at 0.05 (i.e., 95% confidence interval).
@@
2.2 Colour-based Ternary Score
@@
Clearly, the CCIS is not a numerical value but a colour-coded output, like the individual alerts for separate analytes. This clearly distinguishes the CCIS from the normalised immune score.
@@
2.3 Numerical score Derived from the CCIS
We explored an alternative score that would follow the colour-coded scheme, while also producing a numerical value. The intention was to combine the accessibility of the CCIS and the finer precision of the NIS in a single value. As with the latter, the value would range from 0 to 10.
@@ 2.3 Numerical score Derived from the CCIS
Table 2: Mean values and standard deviations found in the NHANES 2003-2016 database for the 13 analytes of the immune scores (same units as above).
@@
Total White Cells (WBC) 6.92 3.12 7.01 2.09
@@
Neutrophils (NEUT) 4 1.85 4.16 1.66
Lymphocytes (LYMPH) 2.11 2.21 2.11 0.79
Monocytes (MONO) 0.57 0.21 0.52 0.18
Eosinophils (EOS) 0.20 0.16 0.17 0.14
Basophils (BAS) 0.04 0.06 0.04 0.05
The behaviour of Fig. 4 illustrates how values at a remove from healthy ones are pushed towards higher values by design, as an alerting mechanism.
@@
  • $w_{a}=\sqrt{100/13}$: the maximum normalised distance for each analyte in the normalised vector of a patient’s readings.
  • $e(g,i)$: the mean (or expected) value for analyte $i$ in group $g$.
  • $\sigma(g,i)$: the standard deviation from the mean.
  • $n(g,i)=(r_{u}(g,i)-r_{l}(g,i))/2$ : the normal (healthy reference) interval length from the mean value (assuming upper and lower limits are equidistant from the mean, an assumption to be revised in the future, both in the NIS and here in the NCCIS).
@@
For global values, the calculation has to take into account both the definition of the CCIS and the fact that each analyte can have proportionally different normal ranges and standard deviations:
@@
2.4 Machine Learning Investigation
The NHANES data on blood analytes (CBC markers) and NIS immune scores across patients were separated by health status for a binary classification problem: healthy vs. unhealthy. The data was preprocessed by imputing the missing NaN values using the mean for each feature. The dataset was then split into training and test sets for classification, model building and evaluation. Test sizes of 0.2 and 0.5 were used with a 10-fold cross-validation for performance metrics analysis. Five statistical machine learning (ML) classifiers were used for initial assessment: Random Forest (RF), Support Vector Machine (SVM), K-Nearest Neighbours (KNN), KMeans clustering, and Gaussian Mixture Model. The latter two clustering methods delivered poorer performances in the data features, thus were removed from the screening. Each model was trained on the training set and evaluated on the held-out test set. Feature importance was analysed for the RF, SVM, and KNN models to identify the most predictive features. Pairwise analysis was also conducted to inspect feature relationships. A correlation analysis was performed to identify any co-linear features before saving the results.
Next, two artificial neural network (ANN) architectures were explored as more flexible artificial intelligence (AI) tools. The ANN models tested on the binary health classification by analytes and immune scores included a feed-forward neural network (FFNN) (also referred to as a multilayer perceptron MLP). The MLP consists of an input layer, one or more hidden layers, and an output layer. The model uses ReLU (Rectified Linear Unit) activation functions for the hidden layers and a sigmoid activation function for the output layer, which is typical for binary classification tasks.
Additionally, in order to make the ANN model more robust, we incorporated a more complex layered architecture with a Transformer network. The Keras and TensorFlow python packages were used for the development of the neural network model. Statistical metrics such as ROC curves (area under the curve- AUC), sensitivity, specificity, mean square error (MSE) and classification accuracy were computed to evaluate the predictive performance of the ML algorithms.
K-Nearest Neighbors (KNN): The K-NN algorithm, implemented using the sklearn.neighbors package, is a non-parametric method used for classification. The primary hyperparameters include: n_neighbors set to a default of 5. The weights parameter determines how the votes of the neighbours are weighted during the decision trees. It was kept at default settings of ’uniform’. The algorithm parameter was also kept in the default state of ’auto’ option, as it attempts to choose the optimal algorithm based on the input data structure. Additionally, the leaf_size parameter was set to a value of 30.
Support Vector Machine (SVM): Available through the sklearn.svm package, this is a powerful supervised learning model. The key hyperparameters include $C$, the regularisation parameter to regulate the trade-off between achieving a low error on the training data and minimising the model’s complexity. $C$ was set to a default value of 1.0. The kernel parameter was kept at ’rbf’, being the default. The gamma parameter (the kernel coefficient for ’rbf’) was kept at a default value ’scale’, which uses $\frac{1}{\text{n\_features}\times\text{X.var()}}$. All other parameters were kept at default.
Random Forest (RF): Implemented using the sklearn.ensemble package, this is an ensemble learning method for classification that constructs multiple decision trees during training and outputs the mode of the classes. The algorithm detects implicit regularisation of patterns. The primary hyperparameters for RF include n_estimators (the number of trees in the forest) at a default value of 100. The criterion parameter that defines the function to measure the quality of a split was kept at ‘gini’. The max_depth was kept at default (None) and min_samples_split was kept at a value of 2.
Artificial Neural Network (ANN - Feedforward Neural Network): A multilayer perceptron with one input layer and two hidden layers was used. The first hidden layer consisted of 64 neurons with ReLU activation, followed by a dropout layer with a dropout rate of 0.5 to prevent overfitting. The second hidden layer had 32 neurons with ReLU activation. The output layer consisted of 1 neuron with Sigmoid activation for binary classification. The primary hyperparameters included the Adam optimiser with a learning rate of 0.001, the loss function (binary cross-entropy), and accuracy metrics. Changing the first hidden layer to 128 neurons and the second hidden layer to 64 neurons did not make a difference as the initial model was already at optimal performance.
Transformer-Based Neural Network (Transformer-ANN): Also implemented using the keras package from tensorflow. One input layer and one output layer (sigmoid activation function) with one neuron each. The input layer was reshaped into a sequence dimension before the attention mechanism. The attention mechanism was followed by dense layers with 64 neurons each and dropout layers. The dropout layer used a dropout rate of 0.5. The self-attention mechanism, using keras.layers.MultiHeadAttention consisted of two additional hyperparameters: The num_heads (Number of attention heads) and the key_dim (size of each attention head) were both set to 2. This was designed to handle sequential data for the detection of long-range dependencies. A flatten layer was applied to flatten the output of the attention layer prior to output (binary classification).
@@ 2.5 NHANES Age Prediction
The sex groups were separated by health status as determined by binning the immune scores and ages by the healthy range thresholds indicated in Table  1, which are as follows: Red Blood Cell (RBC) count (between 3.8 to 5.8), and the following healthy ranges for the multivariate measure: normal Mean Corpuscular Haemoglobin (MCH) (27-33), Haematocrit (HCT) (37-47), lymphocytes (LYMP) (1.0-4.0), and neutrophils (NEUT) (1.8-7.5). Above or below these healthy count ranges, individuals were binned into the unhealthy groups. Further, the error rates analysis for the age prediction between the actual ages and immune-score computed ages was performed using these binned subsets.
The zero counts were removed from the NHANES data prior to all analysis. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table 2 highlights sex-specific differences in immune analyte values, for reference.